OpenStudy (anonymous):

Solve for the variable in the following proportion. n : 1/2 as 6 : 1

OpenStudy (triciaal):

visualize fractions n/(1/2) = 6/1 method one when you divide by a fraction invert and multiply n/(1/2) = n(2/1) now you have 2n = 6 divide by 2 to get n=3 method two multiply across the = n(1) =(1/2)(6) n = 3